Fermi National Laboratory


Accelerator Update for September 5, 2005 - September 7, 2005

Monday, September 5
The day shift began with Operations monitoring store 4370, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations put the Recycler into shot setup at 8:54 AM to stash antiprotons.

The Recycler ended their stash and Pbar resumed stacking at 10:38 AM. MCenter resumed taking beam.

The Recycler began stashing at 2:21 PM.

The stash ended and Pbar resumed stacking and MCenter resumed taking beam at 3:22 PM.

At 7:12 PM, the Booster East Anode Power Supply (EAPS) tripped off. A Booster RF station (BRF8) also tripped off. Operations had to drop the Booster beam intensity with BRF8 off. Operators disconnected BRF8 from the RF system and called in an expert.

Booster beam resumed at 10:23 PM.

Tuesday, September 6
The midnight shift began with Operations monitoring store 4370 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4370 at 8:03 AM.

Operation put the Tevatron (TeV) into shot setup at 8:22 AM.

Water technicians reported at 9:01 AM that there was a LCW leak in the Meson M05 area. They turned off the system to investigate. Meson beam was already off due to shot setup. Later, they found a leak on MT5E-5 and had it repaired by 10:29 AM.

Operations reported at 10:52 AM that power supply techs were beginning work on EAPS. This shut off Booster beam. They also repaired a LCW leak on BRF4.

Operations established store 4371 at 12:03 PM with an initial luminosity of 89.84E30.

Techs completed the Booster EAPS and BRF4 work and Booster beam resumed around 12:25 PM.

MiniBooNE and MCenter resumed taking beam by 12:34 PM. Pbar began a study period.

A Booster expert reported at 4:10 PM that the 400 MeV area chopper was, on occasion, not sending pulsed beam to Booster. He asked another expert to investigate and make adjustments.

TeV experts began a study period around 4:30 PM. The study ended at 4:54 PM.

At 6:11 PM, CDF noted a problem with the antiproton beam. Operators investigated, made some adjustments, and then found that a collimator had moved into the beam. They had the problem fixed by around 7:41 PM.

At 6:45 PM, a Debuncher RF station (DRF1-5) tripped off for the first of three times in about five hours. It reset each time.

Wednesday, September 7
The midnight shift began with Operations monitoring store 4371, with Pbar conducting various stacking studies, and with MiniBooNE and NuMI taking beam.

At 1:19 AM, a Main Injector (MI) RF station (MIRF2) tripped off. It reset okay, but then a few minutes later MIRF11 tripped off. It reset okay, but then ten minutes later MIRF7, 11, and 14 tripped off. They all reset okay.

The Recycler began stashing antiprotons at 2:06 AM.

At 3:14 AM, MIRF2 tripped off. Fifteen minutes later MIRF11 tripped and then two minutes after that MIRF11 and 14 tripped. Everything reset okay.

The Recycler stash ended and Pbar resumed stacking at around 3:15 AM.

At 7:04 AM, DRF1-5 tripped off. It reset okay.

The Recycler began stashing at 7:35 AM.

The Plans for Wednesday
The plans for today are to keep store 4371 in until 4 PM. TeV experts will conduct a two-hour end-of-store study beginning around 1 PM. At 3:30 PM Operations will terminate the store and the Run Coordinator will allow CDF to begin a 16-hour access.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/7/2005   email Fermilab